# Transfer Learning Between U.S. Presidential Elections: How Should We Learn From A 2020 Ad Campaign To Inform 2024 Ad Campaigns?

### v1

@@ 6.1 Setup
We apply our approach to study the effect of running a negative, digital ad campaign against Trump among registered voters in PA for the 2024 U.S. presidential election. The target data is from the PA’s voter registration database as of April 15, 2024, which initially contained $8,716,343$ registered voters. To harmonize with the source data by Aggarwal et al., (2023), we took a subset of voters in the PA database who are between 18 and 55 years old. We also recoded age, political party registration, and voting history in the PA database to match the definitions in the source data. In the end, we had $n_{t}=4,880,729$ registered voters in the target data and the shared covariates ${\bf V}_{i}$ included gender, age groups, party, and a subset of the voting history. The source covariates ${\bf X}_{i}$ included ${\bf V}_{i}$, race, and a richer set of voting history. Figure 6.1 shows all of the covariates. For more details on the data description and data cleaning, see Section E of the Appendix.
For all 67 counties in PA, we estimate the ad effect in Section 6.2. We also conduct a subgroup analysis by gender, urbanicity, and education in Section 6.3. Due to page constraints and since all covariates are discrete, we present the results from the OR estimator and discuss the results from the EIF estimator in Section E of the Appendix; except for some discrepancies noted in Section 7, the two estimators reach the same conclusion. The regression function $\rho_{a}({\bf v})$ is estimated by regressing $\widehat{\mu}_{a}({\bf x})$ on ${\bf v}$. Following Aggarwal et al., (2023), $\mu_{a}({\bf x})$ is estimated by weighted least squares where the weights are the inverse propensity scores. We consider $\gamma_{a}$’s between $-0.05$ and $0.05$ and as discussed above, obtain calibrated sensitivity parameters by partitioning the source data into blue-collar states (i.e., PA, WI, MI) and non blue-collar states (i.e., AZ, NC). Following Aggarwal et al., (2023), $\widehat{\rm CI}_{s_{1}}$ and $\widehat{\rm CI}_{s_{2}}$ in the calibration procedure are based on weighted least squares that regresses the outcome on the treatment and pre-treatment covariates and the weights are the inverse of the propensity scores. Throughout the analysis, we set the significance level at $\alpha=0.05$.
@@
6.2 Ad Effect by Counties: Key Takeaways
When transportability holds, the digital ads against Trump decreases turnout in Fulton county.
@@ 6.2 Ad Effect by Counties: Key Takeaways
When $\gamma_{0}=\gamma_{1}=0$ (i.e., transportability holds), the ad effect is negative and barely significant in Fulton county (95% CI: $[-1.64\%,-0.04\%]$, p-value: $0.04$), and insignificant in all other 66 counties; see Section E of the Appendix for the exact numbers. In other words, if the difference in voter turnout between PA voters in 2024 and the voters in 2020 can be adjusted with ${\bf V}_{i}$, then the ads will be ineffective in almost all counties for the 2024 election, except for Fulton county, which is slightly significant. Interestingly, Fulton county had the largest share of votes for Trump in the 2020 U.S. presidential election, with 85.41% of votes in Fulton county for Trump.
After calibration, (a) Philadelphia county is sensitive for a positive ad effect (i.e., increase turnout), (b) 54 counties are sensitive for a negative ad effect (i.e., decreased turnout), and (c) 12 counties are insensitive for an ad effect.
When $\gamma_{0}\gamma_{1}\neq 0$ (i.e., transportability is violated), the digital ads against Trump can have statistically significant effects among the calibrated values of the sensitivity parameters; see the left panel of Figure 6.2. Philadelphia county is sensitive for a positive effect and 54 counties are sensitive for a negative effect. In words, under a range of unmeasured differences defined by the unmeasured differences between blue-collar and non blue-collar states in the source data, receiving ads against Trump can increase voter turnout in Philadelphia county and decrease voter turnout in 54 counties.
Finally, 12 counties are insensitive for a significant ad effect. In other words, even if the unmeasured differences between the 2020 and 2024 populations exist, but the magnitude of the differences is in the calibration/reference set $\mathcal{C}$ generated from the source data, running ads against Trump will not change voter turnout in 12 counties. We remark that this does not mean that the ads are actually ineffective in the 12 counties; rather, the ads are ineffective if we only consider magnitudes of the unmeasured differences between blue-collar and non blue-collar states in the source data.
@@
In general, we see that the direction of the ad effect after calibration roughly corresponds to Trump’s share of votes in the 2020 U.S. presidential elections; see Figure 6.2. Philadelphia county, which was estimated to have a positive effect, has a history of voting for Democratic presidential candidates by large margins. Also, Bedford, Juniata, and Somerset counties, which were estimated to be sensitive for a negative effect, voted for Trump in 2020 by large numbers; Trump received 83%, 80%, and 77% of the votes from Bedford, Juniata, and Somerset counties, respectively. Combined with the result from Fulton county above, counties that strongly supported Trump in 2020 are likely to have negative effects during the 2024 election.
Figure 6.2: Sensitivity analysis of the ad effect for each county in PA. Left: Counties that are sensitive for a positive effect (i.e., increased turnout; in blue) or a negative effect (i.e., decreased turnout; in red). Counties that are insensitive for an effect are in gray. Right: Trump’s share of votes in the 2020 U.S. presidential election.
@@ 6.3 Subgroup Analysis
After overturning of Roe v. Wade in 2022, many argued that voter turnout will vary substantially by gender and urbanicity, especially compared to past elections (e.g., Shea and Jacobs, (2023); Fair Vote, (2024)). To study whether the effect of the ad will also vary by voter demographics, we estimate the effect of running a negative, digital ad campaign against Trump among 20 subgroups of voters. The 20 subgroups are defined by a three-way interaction between gender (female versus not female), urbanicity (rural versus urban), and education attainment (five levels). Specifically, we use the U.S. Census to obtain information about whether (a) a PA voter lives in a rural or an urban census tract and (b) a PA voter lives in a zipcode with a certain level of educational attainment. Education attainment is categorized by the percentage of people with a Bachelor’s degree or higher and is in increments of 20% (i.e., $(0,20\%],(20,40\%],(40,60\%],(60,80\%],(80,100\%]$). Section E of the Appendix contains further details about the subgroups.
@@
Figure 6.3 summarizes the results. Under transportability, we find some variations in the ad effect among different subgroups of voters, but none of the effects are statistically significant. Voters in urban areas have non-negative ad effects regardless of gender and educational attainment and the magnitude of the effect roughly increases with educational attainment. Among voters in rural areas, the ad effect is positive among females living in areas with high educational attainment and the magnitude of this effect is comparable to voters who live in urban areas. The ad effect is most negative among female voters living in rural areas with low educational attainment. In other words, in the 2024 election, running against Trump can suppress turnout among female voters living in rural areas with low educational attainment.
@@
When transportability is violated and after calibration, the ad effect is sensitive in a negative direction among female voters living in rural areas with moderate to low educational attainment. The ad effect is sensitive in a positive direction among non-female voters living in urban areas with high educational attainment. The ad effect is sensitive in both directions among non-female voters living in a rural area with high educational attainment. In other words, if the unmeasured differences considered in the calibration set $\mathcal{C}$ are plausible, the digital ads against Trump will have a statistically significant effect among six subgroups of PA voters in 2024.
@@
Figure 6.3: Subgroup analysis by gender, urbanicity, and levels of education attainment in a voter’s zipcode. The bar represents the 95% CI of the ad effect under transportability. On the right, statistically significant effects after calibration are colored in blue, red, and orange for positive, negative, and bidirectional effects. Insignificant results are colored gray.
7 Discussion
This paper proposes to use transfer learning with sensitivity analysis to address whether running a digital ad campaign against Trump is effective in changing voter turnout in PA for the 2024 U.S. presidential elections. While not ideal compared to running a randomized trial during the 2024 election, the proposed approach is considerably cheaper as it leverages existing, large-scale experimental data from Aggarwal et al., (2023) and uses sensitivity analysis to account for shifts in context and voter demographics between elections. We present two estimation procedures for the TATE, one based on outcome regression (OR) modeling with bootstrapped CIs (i.e., the recommended procedure) and another based on the EIF. For each procedure, we show that it leads to consistent estimates of the TATE and asymptotically valid $1-\alpha$ CIs; see Section G of the Appendix for a simulation study that corroborates the theoretical conclusions empirically. Finally, inspired by ideas from design sensitivity, we present a calibration procedure based on partitioning the source population and use it to generate a set of reference magnitudes of the sensitivity parameters.
The analyses based on the OR estimator and the EIF estimator were similar, but not identical. Specifically, the subgroup analysis based on the EIF estimator was identical to that in Section 6.3 except the effect is insensitive in the negative direction for non-female voters living in rural areas with more than 80% having Bachelor’s degree or higher. The county-by-county analysis based on the EIF estimator yielded no significant results under transportability. After calibration, the analysis based on the EIF estimator produced three more counties sensitive for a positive effect and 34 fewer counties sensitive for a negative effect than those in Section 6. However, the sensitive counties follow the same trend with respect to percentage of votes for Trump in 2020; see Section E of the Appendix. We remark that the widths of the CIs from the EIF estimator and those from the OR estimator did not uniformly dominate each other.
One major reason for the difference between the EIF estimator and the OR estimator is due to the numerical instability from estimating the density ratio $w({\bf v})$, one of the four nuisance functions in the EIF estimator. The density ratio is difficult to estimate when the positivity condition (i.e., Assumption 2.2) is nearly violated in finite sample and this was the case for less-populated counties in PA. Importantly, an unstable estimate of the density ratio leads to less robust “extrapolation” of the source population into the target population. Given these concerns in our data and the discreteness of $\mathcal{X}$, we decided to present our findings based on the OR estimator.
@@ 7 Discussion
Our framework assumed that the target and source samples are independent and there are no overlapping voters between the two samples. But in our analysis, it’s plausible that a registered voter in PA for the 2020 election remained a registered voter in PA for the 2024 election. Unfortunately, the source data from Aggarwal et al., (2023) does not identify the voter’s residence exactly. Nevertheless, to allay concerns on potentially overlapping voters, we repeated our analysis with a restricted source data consisting of $n_{s}=662,225$ voters from NC and AZ only. The results from this analysis follow the same trends as above, but with less statistically significant results due to a much smaller sample size. Specifically, the county-by-county analysis results in no counties that are significant under transportability. Even after calibration, no counties are sensitive for positive effects and three fewer counties are sensitive for negative effects than those in Section 6.2. Also, the subgroup analysis did not yield any significant effects after calibration. For more details, including various robustness checks related to decisions we made during the data pre-processing step, see Sections E and F of the Appendix.
@@
While restricting the source data to NC and AZ removes concerns on overlapping voters, it increases the implausibility of the transportability since the target population is less similar to the restricted source data (i.e., voters in 2020 in NC and AZ) than the original source data (i.e., voters in 2020 from the five states including PA). We generally believe that the source and the target population should be as similar as possible to minimize bias and thus, we decide to report the analysis where the source data contains voters from five states including PA.
Our framework implicitly assumes that the units in the target and the source data are sampled from an infinite population of voters. But, it may be more appropriate to treat the PA voter registration database as a finite population or a large sample from a finite population. These questions about sampling from a finite population and overlapping units between finite populations raised several interesting, theoretical questions about semiparametric efficiency. Due to space constraints and this paper’s emphasis on application, we address them in an upcoming paper.

### v2

@@ 6.1 Setup
We apply our approach to study the main empirical question from the paper, i.e., what is the effect of running a negative, digital ad campaign against Trump among registered voters in PA for the 2024 U.S. presidential election? The target data is from the PA’s voter registration database as of April 15, 2024, which initially contained $8,716,343$ registered voters. To harmonize with the source data by Aggarwal et al., (2023), we took a subset of voters in the PA database who are between 18 and 55 years old. We also recoded age, political party registration, and voting history in the PA database to match the definitions in the source data. In the end, we had $n_{t}=4,880,729$ registered voters in the target data and the shared covariates ${\bf V}_{i}$ included gender, age groups, party, and a subset of the voting history. The source covariates ${\bf X}_{i}$ included ${\bf V}_{i}$, race, and a richer set of voting history from Aggarwal et al., (2023) and there were $n_{s}=1,999,282$ registered voters from the source data. Figure 6.1 visualizes all of the covariates. For more details on the data description and data cleaning, see Section E of the Appendix.
For all 67 counties of PA, we estimate the ad effect in Section 6.2. We also conduct a subgroup analysis by gender, urbanicity, and education in Section 6.3. Due to page constraints and since all covariates are discrete, we present the results from the OR estimator and discuss the results from the EIF estimator in Section E of the Appendix; except for few discrepancies noted in Section 6.4, the two estimators reach the same conclusion. The regression function $\rho_{a}({\bf v})$ is estimated by regressing $\widehat{\mu}_{a}({\bf x})$ on ${\bf v}$. Following Aggarwal et al., (2023), $\mu_{a}({\bf x})$ is estimated by weighted least squares where the weights are the inverse propensity scores. The density ratio $w^{(k)}({\bf v})$ is estimated by entropy balancing in (4.8). As discussed above, we obtain calibrated sensitivity parameters by partitioning the source data into the rust belt states (i.e., PA, WI, MI) and the sun belt states (i.e., AZ, NC). Following Aggarwal et al., (2023), $\widehat{\rm CI}_{s_{1}}$ and $\widehat{\rm CI}_{s_{2}}$ in the calibration procedure are based on weighted least squares that regresses the outcome on the treatment and pre-treatment covariates and the weights are the inverse of the propensity scores.
Throughout the analysis, the significance level is $\alpha=0.05$. Also, as a reminder, a positive effect means that running negative ads against Trump increased voter turnout whereas a negative effect means that running negative ads decreased voter turnout.
@@
6.2 Ad effect by counties
6.2.1 Ad effect under transportability
@@
When $\Gamma_{0}=\Gamma_{1}=1$ (i.e., transportability holds), the ad effect is negative (i.e., decreased voter turnout when assigned to anti-Trump ads) and barely significant in Fulton county (95% CI: $[-1.64\%,-0.04\%]$, $p=0.04$). In all other 66 counties, the ad effects are insignificant; see panel A of Figure 6.2 for a visual illustration and Section E of the Appendix for the exact numbers. In other words, if the difference in voter turnout between PA voters in 2024 and the voters in 2020 can be completely adjusted with ${\bf V}_{i}$, then the negative ads will be ineffective in almost all counties for the 2024 election, except for Fulton county.
6.2.2 Ad effect with pre-specified $(\Gamma_{0},\Gamma_{1})$s
Next, we study the ad effects for different values of $(\Gamma_{0},\Gamma_{1})$. For brevity, we present two values of $(\Gamma_{0},\Gamma_{1})$ in panel B of Figure 6.2, and defer other values to Section E of the Appendix. We remark that this section mirrors a “traditional” sensitivity analysis discussed in Section 5.1 where the investigator pre-specifies $(\Gamma_{0},\Gamma_{1})$s based on existing, prior beliefs about the unmeasured differences between the 2020 and the 2024 elections.
Suppose $\Gamma_{0}=1.01$ and $\Gamma_{1}=0.99$, i.e., in the control arm, the counterfactual odd of voting in 2024 is $1.01$ times the counterfactual odd in 2020 and in the treated arm, the counterfactual odd of voting in 2024 is $0.99$ times the counterfactual odd in 2020. The ad effect is significant and negative in 51 counties and insignificant in 16 counties. The p-value is the smallest in Fulton ($p=0.021$), followed by Bedford ($p=0.027$) and Juniata ($p=0.042$).
Conversely, suppose $\Gamma_{0}=0.99$ and $\Gamma_{1}=1.01$, i.e., in the control arm, the counterfactual odd of voting in 2024 is $0.99$ times that in 2020 and in the treated arm, the counterfactual odd of voting in 2024 is $1.01$ times that in 2020. The ad effect is significant and positive in Philadelphia county ($p=0.044)$ and insignificant in other counties.
6.2.3 Ad effect with the calibrated set
We use the sensitivity parameters from the calibrated set ${\cal C}$ in Section 5.2 to conduct the sensitivity analysis. With a slight abuse of notation, we use ${\cal C}$ to denote the calibrated set for every county; see the discussion on sampling uncertainty in Section 5.3 and panel C of Figure 6.2 for examples of ${\cal C}$. An illustration of the calibrated results is shown in panel A of Figure 6.3.
Following Definition 5.1, 61 counties are sensitive to transportability within the calibration set. Philadelphia county is sensitive in that its result changed from an insignificant ad effect under transportability to a significant and positive effect when transportability is violated by the amount in the calibrated set ${\cal C}$; we refer to this type of sensitivity as sensitive for a positive effect. Bedford county is sensitive in that its result changed from an insignificant effect under transportability to a significant and negative effect when transportability is violated by the amount in the calibrated set ${\cal C}$; we refer to this type of sensitivity as sensitive for a negative effect. Fulton county is sensitive to transportability in that its result changed from a significant and negative effect under transportability to an insignificant effect when transportability is violated by the amount in the calibrated set ${\cal C}$; we refer to this type of sensitivity as sensitive for an insignificant effect. Overall, 59 counties are sensitive for a positive effect, one county is sensitive for a negative effect, and one county is sensitive for an insignificant effect. The other remaining six counties are insensitive.
In words, the conclusions of the 2024 ad effect can change in 61 counties from their corresponding conclusions under transportability if we consider the magnitudes of unmeasured differences between the sun belt and the rust belt states in 2020. Also, the conclusions of the 2024 ad effect remain unchanged in six counties if we consider the magnitudes of unmeasured differences between the sun belt and the rust belt states in 2020.
Figure 6.2: County-by-county analysis results for 2024 PA voters. Panel A: results under transportability. Panel B: results under two values of $(\Gamma_{0},\Gamma_{1})$. Panel C: an illustration of the analysis pipeline for Philadelphia county and Fulton county. The left column of panel C expands the analysis in panel B for various values of $(\Gamma_{0},\Gamma_{1})$ and each point represents the statistical significance of the ad effect for each $(\Gamma_{0},\Gamma_{1})$; only a few $(\Gamma_{0},\Gamma_{1})$s are displayed for visualization purposes. Gray represents an insignificant effect, blue represents a significant and positive effect, and red represents a significant and negative effect. The middle column of panel C conducts the calibration procedure and the green area is the calibrated set ${\cal C}$. The right column of panel C is the overlap of the two plots and represents the results of the sensitivity analysis with the sensitivity parameters in the calibrated set ${\cal C}$.
Figure 6.3: County-by-county analysis results. Panel A: results of sensitivity analysis under the calibrated set ${\cal C}$. The downward (upward) arrow represents a sensitive effect where the effect changed from an insignificant effect under transportability to a significant and negative (positive) effect. The two gray arrows pointing together represents a change from a significant effect to an insignificant effect. Counties without an arrow are insensitive. Panels B and C: Trump’s share (%) of votes in the 2020 and 2024 U.S. presidential elections, respectively.
6.2.4 Summary of the results and interpretations
From the sensitivity analyses in Section 6.2.2, a small, unmeasured difference between 2020 and 2024 leads to different conclusions about the ad effect in many counties compared to their conclusions under $(\Gamma_{0},\Gamma_{1})=(1,1)$ (i.e., when transportability holds). For example, a small, $0.01$ change in the odds of voting between 2020 and 2024, specifically from $(\Gamma_{0},\Gamma_{1})=(1,1)$ to $(\Gamma_{0},\Gamma_{1})=(1.01,0.99)$, yields many more significant conclusions across counties in PA. Similarly, the sensitivity analysis in Section 6.2.3 based on the calibrated set also suggests that many effects will be sensitive if the odds of voting changed by the amount in the calibrated set. If either the values of the sensitivity parameters in Section 6.2.2 or Section 6.2.3 are plausible, then our paper provides some evidence for the conjecture from Aggarwal et al., (2023) that the ad effect from their 2020 experiment will not generalize to most counties in PA for the 2024 election. Also, based on the direction of the sensitive effects, we believe that with the exception of Philadelphia county, the negative ads against Trump will generally decrease voter turnout in the 2024 election. More generally, because the original effects by Aggarwal et al., (2023) were close to null, we believe our analysis framework is the first to empirically illustrate and validate a simple, but under-appreciated point by Rosenbaum, (2010) that “small effects are sensitive to small [unmeasured] biases” in the context of transfer learning.
Similar to other sensitivity analyses, whether the $(\Gamma_{0},\Gamma_{1})$s considered in Sections 6.2.2 and 6.2.3 are reasonable, unmeasured differences between 2020 and 2024 is at the investigator’s discretion. The two sections provide different ways to interpret the sensitivity analysis and how conclusions of the TATE would change when transportability is violated by a certain amount. We also repeat two cautionary notes from Section 5.3 in that (a) the true unmeasured differences between 2020 and 2024 are not equal to the values of the sensitivity parameters in the two sections, especially the calibrated set from the calibration procedure, and (b) investigators can use different partitions in the calibration procedure to tailor the interpretability of the sensitivity parameter to their specific needs.
@@ 6.2.4 Summary of the results and interpretations
Finally, we emphasize that the direction of the ad effect on voter turnout does not equate to whether the ads will lead to less (or more) votes for Trump. This is because Aggarwal et al., (2023) did not measure information about whom a voter voted for. Nevertheless, we can make well-educated conjectures based on comparing the estimated TATEs in Section 6.2 with the shares of votes for Trump across each county; see Figure 6.3. In general, we see that the direction of the sensitive effect roughly corresponds to Trump’s share of votes in the 2020 and 2024 U.S. presidential elections. Philadelphia county, which was declared to be sensitive for a positive effect, has a history of voting for Democratic presidential candidates by large margins. Also, Bedford, Juniata, and Somerset counties, which were declared to be sensitive for a negative effect, voted for Trump by large numbers; in 2020, Trump received 83%, 80%, and 77% of the votes from Bedford, Juniata, and Somerset counties, respectively. However, we caution readers from over-interpreting this connection as Aggarwal et al., (2023) did not measure which candidate a voter voted for.
@@ 6.3 Subgroup analysis
After overturning of Roe v. Wade in 2022, many argued that voter turnout will vary substantially by gender and urbanicity, especially compared to past elections (e.g., Shea and Jacobs, (2023)). To study whether the ad effect will also vary by voter demographics, we estimate the effect of running a negative, digital ad campaign against Trump among 20 subgroups of voters. The 20 subgroups are defined by a three-way interaction between gender (female versus not female), urbanicity (rural versus urban), and education attainment (five levels). We use the U.S. Census to obtain information about whether (a) a PA voter lives in a rural or an urban census tract and (b) a PA voter lives in a zipcode with a certain level of educational attainment. Education attainment is categorized by the percentage of people with a Bachelor’s degree or higher and is in increments of 20% (i.e., $(0,20\%],(20,40\%],(40,60\%],(60,80\%],(80,100\%]$). Section E of the Appendix contains further details about the subgroups.
@@
Figure 6.4 summarizes the results. Under transportability, we find some variations in the ad effect among different subgroups of voters, but none of the estimated effects are statistically significant. Voters in urban areas have positive ad effects (i.e., increased voter turnout) regardless of gender and educational attainment and the effects roughly increase with educational attainment. Among voters in rural areas, the ad effect is positive among females living in areas with high educational attainment and the magnitude of this effect is comparable to voters who live in urban areas. The ad effect is most negative (i.e., decreased voter turnout) among female voters living in rural areas with low educational attainment.
@@
When transportability is violated and we conduct a sensitivity analysis with the calibrated set ${\cal C}$, the ad effect is sensitive for a negative effect among female voters living in rural areas with moderate to low educational attainment. The ad effect is sensitive for a positive effect among non-female voters living in urban areas with high educational attainment. The ad effect is sensitive in both directions among non-female voters living in a rural area with high educational attainment. Overall, for the unmeasured differences considered in the calibration set $\mathcal{C}$, the digital ads against Trump will be sensitive among 6 of the 20 subgroups of PA voters.
@@
Figure 6.4: Subgroup analysis by gender, urbanicity, and education attainment in a voter’s zipcode. The horizontal bar represents the 95% CI of the ad effect under transportability. The colored boxes represent the results of the sensitivity analysis with the calibrated set ${\cal C}$.
6.4 Summary of diagnostics and robustness checks
We briefly highlight two additional analyses that we conducted to strengthen our conclusions above. A complete list of all the diagnostics and robustness checks can be found in Sections E to G of the Appendix. In particular, Sections E and F of the Appendix discuss robustness checks related to decisions we made during data pre-processing. Section G conducts a simulation analysis on semi-synthetic data.
First, the analyses based on the OR estimator and the EIF-based estimator were similar, but not identical. For example, in Figure 6.5, we see that for all 67 counties, the point estimates between the OR estimator and the EIF-based estimator fall closely to the 45 degree line and all the 95% confidence intervals generated from the two estimators overlap; note that the widths of the CIs from the two estimators did not uniformly dominate one another. Also, the subgroup analysis based on the EIF estimator was identical to that in Section 6.3 under transportability, and yielded a total of eight sensitive effects, one more for a positive, sensitive effect and one more for a negative sensitive effect. Given the simplicity of the OR estimator and the discreteness of $\mathcal{X}$, we decided to present our findings based on the OR estimator.
@@ 6.4 Summary of diagnostics and robustness checks
Second, the statistical theory that underpins our data analysis assumed that the target and source samples are independent and there are no overlapping voters between the two samples. But in our analysis, it’s plausible that a registered voter in PA for the 2020 election remained a registered voter in PA for the 2024 election. Unfortunately, the source data from Aggarwal et al., (2023) does not identify the voter’s residence exactly. Nevertheless, to allay concerns on potentially overlapping voters, we repeated our analysis with a restricted source data consisting of $n_{s}=662,225$ voters from NC and AZ only. The results from this analysis follow the same trends as above, but with less statistically significant results due to a much smaller sample size. Specifically, the county-by-county analysis results in no counties that are significant under transportability. Even after calibration, no counties are sensitive for positive effects and three fewer counties are sensitive for negative effects than those in Section 6.2. Also, the subgroup analysis did not yield any significant effects after calibration.
@@
While restricting the source data to NC and AZ removes concerns about overlapping voters, it makes transportability less plausible since the target population is less similar to the restricted source data than the original source data that includes PA. Since the source and the target population should be as similar as possible to minimize bias, we decided to report the analysis where the source data contains voters from five states including PA.
Figure 6.5: Comparison between the OR and EIF-based estimators for estimating ad effects for every county in PA. In each panel, x- and y- axes represent results from the OR and the EIF-based estimator, respectively. The points represent point estimates and the gray bars represent 95% CIs. The dashed line represents the 45 degree line through origin (i.e., $y=x$).
7 Discussion and future work
This paper proposes a framework to evaluate political ads based on transfer learning with sensitivity analysis and we use the framework to address whether running a digital ad campaign against Trump is effective in changing voter turnout in PA for the 2024 U.S. presidential election. While not ideal compared to running a randomized trial during the 2024 election, the proposed approach is considerably cheaper as it leverages existing, large-scale experimental data from Aggarwal et al., (2023) and uses sensitivity analysis to account for unmeasurable shifts in context and voter demographics between elections. We present two estimation procedures for the TATE, one based on OR modeling with bootstrapped CIs (i.e., the recommended procedure) and another based on the EIF. For each procedure, we show that it leads to consistent estimates of the TATE and asymptotically valid $1-\alpha$ CIs. Finally, inspired by ideas from design sensitivity, we present a calibration procedure based on partitioning the source population and use it to generate a set of reference magnitudes of the sensitivity parameters for the sensitivity analysis.
Beyond elections, our framework provides statistically valid solutions to important, practical issues that arise in transportability and generalizability, such as dealing with mis-matched covariates between the source and the target population, addressing violation of transportability under $\mathcal{X}\neq\mathcal{V}$, providing a theoretical basis for a commonly used bootstrap procedure in transfer learning, and proposing a new calibration procedure without omitting a covariate; see Section 1.2 for a full list of contributions. However, we point out one important problem we did not address in this paper. Our framework assumes that the units in the target and the source data are sampled from an infinite population of voters. But, in some settings, including our election data, it may be more appropriate to treat the source and the target populations as a finite population. These questions about sampling from a finite population raised several interesting, theoretical questions and due to space constraints and this paper’s emphasis on application, we address them in an upcoming paper.